Rx1nyDxyz:

Please help The Grand Canyon is approximately 2 to the 9th power kilometers long. Mariner Valley is a canyon on Mars that is approximately 2 to the 12th power kilometers long. About how many times longer is Mariner Valley than the Grand Canyon? 2 2 to the 3rd power 2 to the 21st power 2 to the 108th power

surjithayer:

\[divide ~2^{12}~by~2^9\] \[\frac{ 2^{12} }{ 2^9 }=?\] \[\frac{ x^a }{ x^b }=x^{a-b}\]

vybez:

2 to the 9th is 512 and 2 to the 12th is 4096 subtract those numbers and you would get 3584 hope this helps 😵